The radius of a circle is increasing at the rate of 0.7 cm/s. What is the rate of increase of its circumference

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
given dr/dt=0.7 cm/s
Rate of increase of circumference=2\pi(dr/dt)
=2(0.7)\pi
=1.4\pi
Thanks and Regards
Shaik Aasif
askIITians faculty